OpenStudy (anonymous):

Laura is baking a cake. The recipe says that she has to mix 64 grams of chocolate powder to the flour. Laura knows that 1 cup of that particular chocolate powder has a mass of 128 grams. She added two over three of a cup of chocolate powder to the flour. Should Laura add more chocolate powder to make the exact recipe, or did she go over and by what amount?

OpenStudy (anonymous):

\[128\div3=42.67\]So, 1/3 cup =42.67\[42.67\times2=85.333\]So, 2/3 cup =85.333 So, she went over by about 21.333 grams.

OpenStudy (anonymous):

So she went over 1/3 cup?

OpenStudy (anonymous):

No, what are your answer choices?

OpenStudy (anonymous):

She went over by 1/3 of a cup She needs to add 1/3 of a cup She needs to add 1/6 of a cup She went over by1/6 of a cup

OpenStudy (anonymous):

She went over by 1/6 cup. \[128/6=21.333\]

OpenStudy (anonymous):

Oh ok, But where did u get 1218/6 from?

OpenStudy (anonymous):

128/6 *

OpenStudy (anonymous):

That is how we find how much chocolate powder is in 1/6 cup.

OpenStudy (anonymous):

Okay. Thanks alot!!

OpenStudy (anonymous):

your welcome.